I need help completing the square for this problem.
16x^2-16x-5=0

So far I have tried this:
16x^2-16x+[1/2(-16)]=5+[1/2(-16)]
16x^2-16x+64=5+64
16(x^2-x+4)=69
16(????????????

I'm lost from this point on. I tried factoring but it's not working for me. I hope you can help me.
Thank you very much
Solve a Quadratic Equation by completing the square.
I was not sure what would be an acceptable answer for your teacher so I worked it all the way out. This problem is a Quadratic Equation. Here is your solution.

Answer by Msgenes(10)

Solve a Quadratic Equation by completing the square.
I was not sure what would be an acceptable answer for your teacher so I worked it all the way out. This problem is a Quadratic Equation. Here is your solution.

16x^2-16x-5=0
(16x^2-16-5)/16=0
x^2-1x-5/16=0
x^2-1x-5/16+5/16=0+5/16
x^2-1x=5/16
( -1/2)^2=1/4
x^2-1x+1/4=5/16+1/4(4/4)
(x-1/2)^2=5/16+4/16
(x-1/2)^2=9/16
Sqrt(x-1/2)^2 )= Sqrt(9/16)
x-1/2=3/4
x=1/2±3/4(If this is not your final solution according to your teacher, finished solution continues below)
x-1/2+1/2=3/4+1/2(2/2)
x=3/4+2/4
x=5/4
x-1/2=-3/4
x-1/2+1/2=1/2(2/2)-3/4
x=2/4-3/4
x=-1/4
Good Luck. :)